# Light Rings, Accretion Disks and Shadows of Hayward Boson Stars in asymptotically AdS Spacetime

### v1

@@ 4.4 Accretion Disk and Shadows
Figure 15: $\tilde{r}_{\mathrm{ISCO}}$, $L_{\mathrm{ISCO}}$, and $E_{\mathrm{ISCO}}$ versus $\tilde{\omega}$ for $\tilde{q}=0.56$. The minimum of $\tilde{\omega}$ is $0.0001$.
@@
Let $E_{\mathrm{em}}$ denote the energy of photons emitted from the accretion disk, and $E_{\mathrm{obs}}$ denote the energy received by the observer. Incorporating both gravitational redshift and Doppler effects, the redshift factor can be expressed as Luminet1979
@@
Using the fixed observation frequencies of the Event Horizon Telescope (EHT) (1.3 mm AkiyamaAlberdiAlef2019e or 0.87 mm RaymondDoelemanAsada2024 ), we can invert the emission spectrum distribution of the accretion disk through Equation (33). Defining $\lambda_{0}$ as the observed wavelength and $\lambda$ as the emission wavelength from the disk, they satisfy:
@@
The first row in Fig. 17 shows imaging results at different inclination angles for $\tilde{\omega}=0.1$. Fig. 17a from outside to inside is divided into three discontinuous regions, they are correspond to primary image, secondary image I, and the secondary image II, respectively. Where the primary image and secondary image II correspond to the upper side of the disk, while the secondary image I correspond to the lower side of the disk, respectively. This phenomenon of multiple secondary images appearing in the shadow region has also been observed recently in other studies GyulchevNedkovaVetsov2021 ; Tsupko2022 ; OlmoRubiera-GarciaSaez-ChillonGomez2022 ; Rosa2023 ; RosaMacedoRubiera-Garcia2023 ; Gao2025 . Based on this characteristic, the observed target can be ruled out as a Schwarzschild black hole.
To elucidate the image configuration of the Fig. 17a, Fig. 18 illustrates the correspondence between photon trajectories and spectral images. In Fig. 18a: Ray 1 corresponds to the primary image, in which the position vector of photons (relative to the boson star center) rotates $\pi/2$ from the observer to the disk; Ray 2 correspond to the secondary image I, in which the position vector of photons rotates $23\pi/$ 2; Ray 3 correspond to the secondary image II, in which the position vector of photons rotates $25\pi/$ 2 and its trajectory exhibits double inflection points.
The spectral distribution of secondary image II is different from that of secondary image I. In order to see that clearly, we present an enlarged view of the Fig. 18b in Fig. 19. From Fig. 19a, we observe that the wavelength increases gradually from the inside to the outside on the secondary image, a distribution consistent with that of the primary image. In contrast, as shown in Fig. 19b, the wavelength on the secondary image II decreases from the inside to the outside.
To better understand the spectral distribution in Fig. 19b, we show in Fig. 20 the relationship between the intersection points of photon trajectories with the accretion disk and the corresponding incident angles. Here, the incident angle refers to the angle between the light from the observer and the line connecting the observer and the center of the boson star. It can be seen that as the incident angle decreases, the intersection points move progressively from the inner to the outer regions of the disk. This results in a corresponding decrease in observed wavelength from the inside to the outside in the spectrum.
@@
Figure 19: Magnified view of the central region in Figure 18b, with (a) fov = $5^{\circ}$, and (b) fov = $1^{\circ}$ (only secondary image II).

### v2

@@ IV.4 Accretion Disk and Shadows
Figure 15: $\tilde{r}_{\mathrm{ISCO}}$, $L_{\mathrm{ISCO}}$, and $E_{\mathrm{ISCO}}$ versus $\tilde{\omega}$ for $\tilde{q}=0.56$. The minimum of $\tilde{\omega}$ is 0.00010.00010.0001.
@@
Let $E_{\mathrm{em}}$ denote the energy of photons emitted from the accretion disk, and $E_{\mathrm{obs}}$ denote the energy received by the observer. Incorporating both gravitational redshift and Doppler effects, the redshift factor can be expressed as [75]
@@
Using the fixed observation frequencies of the Event Horizon Telescope (EHT) (1.3 mm [78] or 0.87 mm [79]), we can invert the emission spectrum distribution of the accretion disk through Equation (33). Defining $\lambda_{0}$ as the observed wavelength and $\lambda$ as the emission wavelength from the disk, they satisfy:
@@
The first row in Fig. 17 shows imaging results at different inclination angles for $\tilde{\omega}=0.1$. Fig. 17(a) from outside to inside is divided into three discontinuous regions, they are correspond to primary image, secondary image I, and the secondary image II, respectively. Where the primary image and secondary image II correspond to the upper side of the disk, while the secondary image I correspond to the lower side of the disk, respectively. This phenomenon of multiple secondary images appearing in the shadow region has also been observed recently in other studies [80, 81, 82, 83, 49, 84]. Based on this characteristic, the observed target can be ruled out as a Schwarzschild black hole.
To elucidate the image configuration of the Fig. 17(a), Fig. 18 illustrates the correspondence between photon trajectories and spectral images. In Fig. 18(a): Ray 1 corresponds to the primary image, in which the position vector of photons (relative to the boson star center) rotates $\pi/2$ from the observer to the disk; Ray 2 correspond to the secondary image I, in which the position vector of photons rotates $3\pi/2$; Ray 3 correspond to the secondary image II, in which the position vector of photons rotates $5\pi/2$ and its trajectory exhibits double inflection points.
The spectral distribution of secondary image II is different from that of secondary image I. In order to see that clearly, we present an enlarged view of the Fig. 18(b) in Fig. 19. From Fig. 19(a), we observe that the wavelength increases gradually from the inside to the outside on the secondary image, a distribution consistent with that of the primary image. In contrast, as shown in Fig. 19(b), the wavelength on the secondary image II decreases from the inside to the outside.
To better understand the spectral distribution in Fig. 19(b), we show in Fig. 20 the relationship between the intersection points of photon trajectories with the accretion disk and the corresponding incident angles. Here, the incident angle refers to the angle between the light from the observer and the line connecting the observer and the center of the boson star. It can be seen that as the incident angle decreases, the intersection points move progressively from the inner to the outer regions of the disk. This results in a corresponding decrease in observed wavelength from the inside to the outside in the spectrum.
@@
Figure 19: Magnified view of the central region in Fig. 18(b), with (a) fov = $5^{\circ}$, and (b) fov = $1^{\circ}$ (only secondary image II).